EVERYBODY IN A A ROOM SHAKES HANDS WITH EVERY BODY ELSE.THE TOTAL NUMBER OF HANDSHAKES IS 66.THE TOTAL NUMBER OF PERSONS IN THE ROOM IS?

a)11 b)12 c)13 d)14

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
total no of handshakes = nC2 = 66
n (n-1) /2 = 66
ior n(n-1) = 66*2 = 11*12
No. of people = 12
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty